Ahmned wants to buy a motorcycle that costs today $30,000 dollars His monthly net salary is $4000 and his monthly expenses are $3300. The company selling the motorcycle offers loans as presented in table below:
#of monthly payments % Interest p.a. comp, monthly
Alternative 1 $36 5.50%
Alternative 2 $48 6.52%
Alternative 3 $60 6.55%

What is the best economic alternative for Ahmed having into account his income and expenses?
1
Expert's answer
2018-09-03T15:28:09-0400
Ahmned wants to buy a motorcycle that costs today $30,000 dollars His monthly net salary is $4000 and his monthly expenses are $3300, so he can pay monthly payment of 4000 - 3300 = $700.
Alternative 1: n = 36, i = 5.50%, monthly payment MP = 30,000*(1 + 0.055/12)^36/36 = $982.46 > $700.
Alternative 2 : n = 48, i = 6.52%, monthly payment MP = 30,000*(1 + 0.0652/12)^48/48 = $810.66 > $700.
Alternative 3: n = 60, i = 6.55%, monthly payment MP = 30,000*(1 + 0.0655/12)^60/60 = $693.13 < $700, so it is the best economic alternative for Ahmed having into account his income and expenses.